shift summary for DAY 7/21/99

Arun Tripathi (arun@physics.ucla.edu)
Wed, 21 Jul 1999 14:03:39 -0700 (PDT)

Shift Summary for DAY shift, 21 July 99

On Shift: E. Swallow, A. Tripathi

Hours of beam delivered at all intensities: 0.0
Hours of beam delivered at nominal intensity: 0.0
Hours of E832 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 1.0

Accounting of lost time:
Accelerator maint.

========================================================================
Runs ended this shift:
Run 13956 - LASER SCAN
Run 13962 - PED
Tapes brought to FCC this shift: None

========================================================================
Accesses during shift: Several, see below

========================================================================
Overview of shift:

No beam, TeV down.
A bunch of things were fixed

========================================================================
Description of Shift:

7:19 Start a long laser scan.
Mounted onsplit tape QKLA09 in drive tps2d3 on KTeV3
7:30 Earl ramped down NM4AN to low residual field. Finds -1.1G
field after ramping down. Used to be ~0.5G in previous run.
8:15 Vacuum people show up and make an access to change the ion
gauge and add protection to DP1.

10:00 Angela and Arun make an access to collect DC gas samples.

10:50 Marj and Patty made an access to swap the preamp and postamp
for DC4YU0. Installed the cards that had been tested by Miguel.
12:05 Elliott swapped azlat in crate 53, slot 13. All bits now work.

12:15 Z-monitor now works.

13:02 Elizabeth and 0.5*(Patty+Arun) make an access to swap BA channel
77.

13:30 Bud installed a raditor on the air intake for the alcohol bubbler
air intake. Works great now. Temp -5.1 C.

15:00 ECS "tweaked" EPICURE/RONALD scripts to make sure residual fields
are below 0.5 Gauss. More details in logbook.

NM4AN left with low residual field and NEG polarity.
When beam comes, it is ready for MUON run then AN832UP for data.


========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Note that the tapes half sticking out from the drives now are only half
full. For the next data run, APPEND to these tapes.